OpenStudy (anonymous):

The product of two negative consecutive integers is 5 more than their sum. Find the integers.

OpenStudy (anonymous):

-2 and -1

OpenStudy (anonymous):

can you show me the formula?

OpenStudy (anonymous):

let the numbers be -x and -x=1 Now, with the given condition \[-x * (-x+1) = -x + (-x+1) + 5\] \[x ^{2} - x = x - x +1 +5\] \[x ^{2} - x = 6\] \[x ^{2} - x - 6 = 0\] Now calculate

OpenStudy (anonymous):

suppose two consecutive integers are -x and (-x+1) now according to problem (-x)(-x+1) = (-x)+(-x+1) + 5 x^2 -x - 6= 0 so you will got two root for that x=-3 and x= 2 So the value of -x is 3 or -2 but it is negative integer so Integers will be -2 and -1